The following data relate to labor cost for production of 3,300 cellular telephones:

Actual: 2,230 hrs. at $14.90
Standard: 2,190 hrs. at $15.10

a. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Rate variance $ Favorable
Time variance $ Unfavorable
Total direct labor cost variance $ Unfavorable
Answer #1

Labor rate variance = (SR-AR) *AH

= (15.40-14.90)*2230

= -446 Favourable

Labor time variance = (SH-AH) * SR

= (2190-2230)*15.40

= 604 Unfavorable

Labor cost variance = 446F + 604U

= 158 Unfavorable
Answer #1

Labor rate variance = (SR-AR) *AH

= (15.40-14.90)*2230

= -446 Favourable

Labor time variance = (SH-AH) * SR

= (2190-2230)*15.40

= 604 Unfavorable

Labor cost variance = 446F + 604U

= 158 Unfavorable